Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

What is the probability of flipping a fair coin three

Expert replies
by VJesus12 » Thu Mar 01, 2018 6:02 am
What is the probability of flipping a fair coin three times and the coin landing on heads on exactly two flips?

A. 3/8
B. 5/8
C. 7/8
D. 1/8
E. 1/4

The OA is the option A.

Experts, what are the formulas that I should use here? Can you help me, please? I don't know how to start?
Join the discussion
Source: — Problem Solving |

by [email protected] » Thu Mar 01, 2018 10:33 am
Hi VJesus12,

We're asked for the probability of flipping a fair coin three times and the coin landing on heads on exactly two flips. Since the number of possible outcomes is so small, we can solve this question by simply listing out the possibilities and answering the exact question that is asked.

Since we're flipping a coin 3 times, there are only (2)(2)(2) = 8 possible outcomes. They are...

HHH
HHT
HTH
THH

TTT
TTH
THT
HTT

We're asked for the probability of getting exactly 2 HEADS from those 3 tosses. There are three ways (out of 8 total) to get exactly 2 heads (HHT, HTH and THH).

Final Answer: A

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Scott@TargetTestPrep » Fri Mar 02, 2018 10:05 am
VJesus12 wrote:What is the probability of flipping a fair coin three times and the coin landing on heads on exactly two flips?

A. 3/8
B. 5/8
C. 7/8
D. 1/8
E. 1/4
We can assume the first 2 flips are heads (H) and the last flip is tails (T). Thus:

P(H-H-T) = 1/2 x 1/2 x 1/2 = 1/8

However, we need to determine in how many ways we can get 2 heads and 1 tail. That number will be equivalent to how many ways we can organize the letters H-H-T.

We use the indistinguishable permutations formula to determine the number of ways to arrange H-H-T: 3!/2! = 3 ways. (Note: The 3 ways are H-H-T, H-T-H, and T-H-H.)

Each of these 3 ways has the same probability of occurring. Thus, the total probability is:

1/8 x 3 = 3/8

Answer: A

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

by BTGmoderatorRO » Fri Mar 02, 2018 11:43 am
A coin has two parts; head(H) and a tail(T). Flipping a coin will give 2 outcome H or T
Let's assume the first flip of the coin gives us T and the last two flip gives us H.
therefore, Pr(T-H-H)= 1/2 * 1/2 * 1/2 = 1/8
So, let's find the number of times we can have two head H and a tail T. This can only be done on the number of possible ways we can arrange the T-H-H.
We have T-H-H, H-T-H, H-H-T = 3ways
Each ways has same probabilty of 1/3.
Therefore the total probaility= 1/3 * 1/3 * 1/3 = 3/8
Hence the correct answer is a
Join the discussion